# Geometric, topological and dynamical properties of conformally symplectic systems, normally hyperbolic invariant manifolds, and scattering maps

### v1

@@ D.2. The Gaussian thermostat
Several models of non-equilibrium thermodynamics are based on introducing some forcing as well as some dissipation to keep energy constant. See [WL98].
These systems have several extra structures, but they lead to (see [WL98, equation (2.4)]) maps that satisfy:
@@
The form $\omega$ and the factor $\eta(x)$ used in [WL98] satisfy several other properties that lead to other consequences, but they will not be used below.
@@ D.3. Vanishing lemmas for the generalized models
In the case of the model (D.2), which corresponds to the presymplectic case studied in 2.10, we can take171717 Since $\eta_{\pm}$ are just bounds, one can use sharper bounds to improve the conditions of theorems. Some simple improved bounds used in [WL98] are, for any $K$: $\eta_{-}=\inf_{x}\left(|\eta(x)\eta(f(x))\cdots\eta(f^{K}(x))|\right)^{1/(K+1)}$, $\eta_{+}=\sup_{x}\left(|\eta(x)\eta(f^{-1}(x))\cdots\eta(f^{-L}(x))|\right)^{1/(K+1)}$. $\eta_{+}=\sup_{x}|\eta(x)|$, $\eta_{-}=\inf_{x}|\eta(x)|$.
@@
D.4. Geometric properties of the scattering map
@@
For both models in Section D.1 and  D.2 we can use the proofs of simplecticity of scattering map in Sections 8.2.5, 8.2.7, using the appropriate assumptions on rates. As remarked in the text, these proofs do not require that $\omega$ is closed or non-degenerate. So, they can be used modulo changing the assumptions181818 As it turns out, the proof in Section 8.2.6 can also be adapted. Taking $d$ of (D.2), we obtain that $f^{*}d\omega=\eta d\omega+d\eta\wedge\omega$. The geometric setup of [WL98] implies that $d\eta\wedge\omega=0$. on the rates to include $\eta_{\pm}$ instead of $\eta$.

### v2

@@ D.2. The Gaussian thermostat
Several models of non-equilibrium thermodynamics are based on introducing some forcing as well as some dissipation to keep energy constant. See [wojtkowski1998conformally].
These systems have several extra structures, but they lead to (see [wojtkowski1998conformally, equation (2.4)]) maps that satisfy:
@@
The form $\omega$ and the factor $\eta(x)$ used in [wojtkowski1998conformally] satisfy several other properties that lead to other consequences, but they will not be used below.
@@ D.3. Vanishing lemmas for the generalized models
In the case of the model (D.2), which corresponds to the presymplectic case studied in 2.10, we can take [^20] Since $\eta_{\pm}$ are just bounds, one can use sharper bounds to improve the conditions of theorems. Some simple improved bounds used in [wojtkowski1998conformally] are, for any $K$: $\eta_{-}=\inf_{x}\left(|\eta(x)\eta(f(x))\cdots\eta(f^{K}(x))|\right)^{1/(K+1)}$, $\eta_{+}=\sup_{x}\left(|\eta(x)\eta(f^{-1}(x))\cdots\eta(f^{-L}(x))|\right)^{1/(K+1)}$. $\eta_{+}=\sup_{x}|\eta(x)|$, $\eta_{-}=\inf_{x}|\eta(x)|$.
@@
D.4. Geometric properties of the scattering map for the generalized models
@@
For both models in Section D.1 and  D.2 we can use the proofs of simplecticity of scattering map in Sections 8.2.5, 8.2.7, using the appropriate assumptions on rates. As remarked in the text, these proofs do not require that $\omega$ is closed or non-degenerate. So, they can be used modulo changing the assumptions [^21] As it turns out, the proof in Section 8.2.6 can also be adapted. Taking $d$ of (D.2), we obtain that $f^{*}d\omega=\eta d\omega+d\eta\wedge\omega$. The geometric setup of [wojtkowski1998conformally] implies that $d\eta\wedge\omega=0$. on the rates to include $\eta_{\pm}$ instead of $\eta$.